OpenStudy (anonymous):

the indianapolis police dept employed 1,000 officers and made 37,493 arrests. at this rate , how many arrests could be made if the numbers of officers were increased to 2,500

OpenStudy (anonymous):

37493/1000 = x/2500 for your ratio x = (2500) x (37493/1000)

OpenStudy (anonymous):

is the answer 93732.5

OpenStudy (anonymous):

Excellent! Good job!

OpenStudy (anonymous):

Is this making sense to you?

OpenStudy (anonymous):

yes thank you